OpenStudy (anonymous):

-3square root 18 -2square root 18 +3square root 2

OpenStudy (anonymous):

write \[\sqrt{18}=3\sqrt{2}\] and continue

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow -3\sqrt{18} = -3\sqrt{2 * 3 * 3} = -3*3\sqrt2 = -9\sqrt{2} }\) Just get them like this and get a common radical.

OpenStudy (anonymous):

thanks